# Linear cost mutual information estimation and independence test of similar performance as HSIC

### v1

@@ II-E Choice of basis $B$
Examples in this article use $m=4$ up to kurtosis with $m^{2}=16$ features, which through cross-validation usually turn out optimal e.g. for financial data as in Fig. 2, can be optimized for some specific data this way, or e.g. trying to recognize small rotations for like in Fig. 1.
@@
Dependence between samples should be usually detectable through such pairwise dependencies. However, in theory there can be only more subtle higher order dependencies - HCR independence test would need to include to be able to detect, evaluate, describe. For example by adding to $B_{x}$ and/or $B_{y}$ basis $\sum_{i}\textrm{sign}(j_{i})=2$ indexes, e.g. $a_{11,1}$ would include triplewise dependence like in Fig. 3: that with change of both coordinates, there is change of expected value in the third one. The cost is larger number of features: $O(d^{2})$ for pairwise, $O(d^{3})$ for triplewise, and so on.
We can also perform some preprocessing of data to try to emphasize dependencies, hopefully making it more likely for low order e.g. pairwise test to find them, like PCA (principal component analysis) - rotation of coordinates to eigenvectors of correlation matrix, separately for both samples. Also e.g. using SVD (singular value decomposition) of $a_{\textbf{k},\textbf{j}}$ as $|B^{+}_{x}|\times|B_{y}^{+}|$ matrix, we can optimize both used bases. We can freely perform such optimization working on a family of data samples of similar behavior. However, performing such optimization separately on a single tested sample would also amplify its noise - it would need modification of independence test to include it, e.g. based on random matrix theory, or synthetic empirical simulations.
II-F HCR independence test used in benchmarks
@@
Accepting $H_{0}$ hypothesis if returned Pr(independent) is above some chosen significance level, or reject otherwise.
@@
It is also worth to consider different final tests if all $M$ normalized features are from $N(0,1)$ distribution - testing min and max is very sensitive to dependencies, but could lead to false positives, and ignores intermediate values. It can be extended from $2$ to $2k$ extreme values by sorting $\hat{a}$ to $\textrm{Pr}(\textrm{independent})\approx$
@@
Where $2^{2k}$ is rescaling to get approximately 1 for independent. To include all values we could use e.g. Anderson-Darling, Shapiro-Wilk tests, $\chi^{2}$, permutation, or log-likelihood test.
II-G Toward kernel methods: global HCR vs local HSIC basis
@@
  • global basis often allows to describe density with a reasonable number of features (moments), allowing more practical description e.g. fo,x’)r the discussed reduced complexity,
@@
  • above HCR kernel matrices include all $B_{x}\times B_{y}$ features, while for HSIC Gaussian kernel only single $\sigma$ is used:
@@ III-E HSIC as a stastical test
IV Summary and further work
There was proposed application of HCR mutual information evaluation for independence test, and comparison with state-of-art HSIC. In much lower complexity, finally practical also to find subtle dependencies between very large data samples, it usually obtains even better sensitivity, additionally providing evaluation of mutual information, and interpretable description of the found dependencies, allowing to model joint distribution between data samples. We can further predict conditional probabilities - e.g. using HCR techniques from [14, 15].
As further work, the proposed tests likely can be improved, also from theoretical perspectives, choice of details especially of the final normality test - improving sensitivity and reducing false positive probability, e.g. Fig. 5 uses some standard ones.
There is also mentioned promising research direction of basis optimization e.g. to make it more likely for pairwise test to find dependencies - like initial PCA of data samples, or SVD of $a_{\textbf{j},\textbf{k}}$ matrix to optimize basis. However, it also amplifies noise - should be used on a family of similar samples to optimize basis for the given task, or for single sample using modified independence test e.g. based on random matrix theory.
We should also search for applications, like finding and exploiting subtle dependencies, or information bottleneck neural network training. Also compare with state-of-art mutual information estimation like [9].

### v2

@@ II-F Choice of basis $B$
Examples in this article use $m=4$ up to kurtosis with $m^{2}=16$ features, which through cross-validation usually turn out optimal e.g. for financial data as in Fig. 2, can be optimized for some specific data this way, or e.g. trying to recognize small rotations like in Fig. 1.
@@
Dependence between samples should be usually detectable through such pairwise dependencies. However, in theory there can be only more subtle higher order dependencies - HCR independence test would need to consciously include to be able to detect, evaluate, describe. For example by adding to $B_{x}$ and/or $B_{y}$ basis $\sum_{i}\textrm{sign}(j_{i})=2$ indexes, e.g. $a_{11,1}$ would include triplewise dependence like in Fig. 3: that with change of two coordinates, there is change of expected value of the third one. The cost is larger number of features to estimate and analyze: $O(d^{2})$ for pairwise, $O(d^{3})$ for triplewise, and so on.
II-G PCA, CCA, SVD optimizations
We can also perform some preprocessing of data to try to emphasize dependencies, hopefully making it more likely for low order e.g. pairwise test to find them, like PCA (principal component analysis) - rotation of coordinates to eigenvectors of correlation matrix, separately for both samples.
PCA optimization often amplifies noise, e.g. Canonical Correlation analysis (CCA) allows to repair it - optimizing basis to maximize correlations, as discussed for HCR in [14]. For independence test we could use it to maximize correlations between $f_{\textbf{j}}(\textbf{x})$ and $f_{\textbf{k}}(\textbf{y})$ values in bases for both samples.
For final $a_{\textbf{j},\textbf{k}}$ as $|B^{+}_{x}|\times|B_{y}^{+}|$ matrix we can also use SVD (singular value decomposition) into $U\Sigma V^{T}$ for orthogonal $U^{T}U=V^{T}V=I$ and diagonal $\Sigma$ of singular values, like in Fig. 5 we can optimize both used bases $f_{B_{x}}(\textbf{x})=(f_{\textbf{j}}(\textbf{x}))_{\textbf{j}\in B_{x}}$, $f_{B_{y}}(\textbf{x})=(f_{\textbf{k}}(\textbf{y}))_{\textbf{k}\in B_{y}}$ multiplying them by $U$ and $V$, to represent joint density:
$\rho(\textbf{x},\textbf{y})=(f_{B_{x}}(\textbf{x}))(a_{\textbf{j},\textbf{k}})(f_{B_{y}}(\textbf{y}))^{T}=\left((f_{B_{x}}(\textbf{x}))U\right)\Sigma\left((f_{B_{y}}(\textbf{y}))V\right)^{T}$ (18)
We can freely perform such optimizations working on a family of data samples of similar behavior, using PCA, CCA or SVD of e.g. averaged samples, matrices. However, performing such optimization separately on a single tested sample would also amplify its noise - like in Fig. 5, it needs modification of independence test to include it, e.g. based on random matrix theory, or synthetic empirical simulations.
II-H HCR independence test used in benchmarks
@@
Accepting $H_{0}$ hypothesis if returned Pr(independent) is above some chosen significance level $\alpha$, or reject otherwise.
@@
It is also worth to consider different final tests if all $M$ normalized features are from $N(0,1)$ distribution - testing min and max is very sensitive to dependencies, but could lead to false positives, and ignores intermediate values. As we can see in Fig. 5, also non-extreme values can be valuable to distinguish signal from noise, suggesting to extend independence test to even more sensitive: by using $2k\leq M$ extreme values, $i$-th sorted $\hat{a}$ as $\textrm{sort}(\hat{a})_{i}$ for $\textrm{Pr}(\textrm{independent})\approx$
@@
Where $2^{2k}$ is rescaling to get approximately 1 for independent. Analogously, as in Fig. 5, we can perform such independence test after SVD basis optimization, comparing singular values of $\hat{a}_{\textbf{j},\textbf{k}}$ instead of the original coefficients.
To include all the values we could also use e.g. Anderson-Darling, Shapiro-Wilk, $\chi^{2}$, permutation, or log-likelihood test.
II-I Toward kernel methods: global HCR vs local HSIC basis
@@
  • global basis often allows to describe density with a reasonable number of features (moments), allowing more practical description e.g. for the discussed reduced complexity,
@@
  • above HCR kernel matrices include all $|B_{x}^{+}||B_{y}^{+}|$ features, while for HSIC Gaussian kernel only single $\sigma$ is used:
@@
maybe worth to extend e.g. to multiple $\sigma_{i}$: $\sum_{i}\exp\left(-\frac{\|\textbf{x}-\textbf{x}^{\prime}\|_{2}}{2\sigma_{i}^{2}}\right)$.
@@ III-E HSIC as a stastical test
IV Summary and further work
There was proposed application of HCR mutual information evaluation for independence test, and comparison with state-of-art HSIC. In much lower complexity, finally practical to also work with very large data samples - crucial to find very subtle dependencies, it usually obtains even better sensitivity. Additionally, HCR provides evaluation of mutual information, and interpretable description of the found dependencies, allowing to model joint distribution between data samples, including significance evaluation of contributions. Finding dependencies, We can further exploit them to predict conditional densities - e.g. using HCR techniques from [17, 18, 14].
As further work, the proposed tests likely can be improved, also from theoretical perspectives, choice of details especially of the final normality test - improving sensitivity and reducing false positive probability, e.g. Fig. 6 uses some standard ones. The proposed test (19) of sorting and individually comparing multiple values with distributions of sorted (15) e.g. $N(0,1)$ seems promising general approach for normality test, also for other distributions - worth to be explored.
There is also mentioned promising research direction of basis optimization e.g. to make it more likely for pairwise test to find dependencies - like initial PCA or CCA of data samples, or directly SVD of $a_{\textbf{j},\textbf{k}}$ matrix to optimize basis as in Fig. 5. It also amplifies noise - should be used on a family of similar samples to optimize basis for the given task, or for single sample using modified independence test e.g. based on random matrix theory or simulations randomly generating matrices.
Finally we should explore applications, like finding and exploiting subtle dependencies e.g. in financial data, or information bottleneck neural network training. Also compare with state-of-art mutual information estimation like [9].